Consider the following two mutually exclusive projects:

Year Cash Flow (A) Cash Flow (B)
0 –$ 364,000 –$ 52,000
1 46,000 25,000
2 68,000 22,000
3 68,000 21,500
4 458,000 17,500


Whichever project you choose, if any, you require a return of 11 percent on your investment.

a-1. What is the payback period for each project? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
  



a-2. If you apply the payback criterion, which investment will you choose?
  

Project A

Project B



b-1. What is the discounted payback period for each project? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
  


b-2. If you apply the discounted payback criterion, which investment will you choose?
  

Project A

Project B



c-1. What is the NPV for each project? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
  


c-2. If you apply the NPV criterion, which investment will you choose?
  

Project A

Project B



d-1. What is the IRR for each project? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
  


d-2. If you apply the IRR criterion, which investment will you choose?
  

Project A

Project B



e-1. What is the profitability index for each project? (Do not round intermediate calculations and round your answers to 3 decimal places, e.g., 32.161.)
  


e-2. If you apply the profitability index criterion, which investment will you choose?
  

Project A

Project B



f. Based on your answers in (a) through (e), which project will you finally choose?
  

Project A

Project B
Answer #1

a-1. What is the payback period for each project?

Project A

Year 0 1 2 3 4
Cashflow(in $)          (364,000)              46,000                     68,000                 68,000               458,000
Cumulative Cashflow(in $)          (364,000)          (318,000)                (250,000)            (182,000)               276,000

Payback Period = A+(B/C)

where

A - last period containing negative cumulative cash flow

B - absolute value of cumulative cash flow in A

C - cash flow during the period after A

Payback Period = 3 + (182000/458000)

= 3 + 0.39737991266

= 3.40 Years

Project B

Year 0 1 2 3 4
Cashflow(in $)             (52,000)              25,000                     22,000                 21,500                 17,500
Cumulative Cashflow(in $)             (52,000)            (27,000)                     (5,000)                 16,500                 34,000

Payback period = 2 + (5000/21500)

= 2 + 0.23255813953

= 2.23 years

a-2. If you apply the payback criterion, which investment will you choose?

Project B as it has lower Payback period.

b-1. What is the discounted payback period for each project?

Project A

Year 0 1 2 3 4
Cashflow(in $)          (364,000)              46,000                     68,000                 68,000               458,000
PVF @11%                          1                 0.901                       0.812                   0.731                   0.659
Discounted Cashflow (Cash flow * PVF)          (364,000)              41,441                     55,190                 49,721               301,699
Cumulative Cashflow(in $)          (364,000)          (322,559)                (267,368)            (217,647)                 84,052

Discounted Payback Period = 3 + (217647/301699)

= 3 + 0.72140444615

= 3.72 Years

Project B

Year 0 1 2 3 4
Cashflow(in $)             (52,000)              25,000                     22,000                 21,500                 17,500
PVF @11%                          1                 0.901                       0.812                   0.731                   0.659
Discounted Cashflow (Cash flow * PVF)       (52,000)        22,523               17,856           15,721           11,528
Cumulative Cashflow(in $)             (52,000)            (29,477)                  (11,622)                   4,099                 15,627

Discounted payback period = 2+(11622/15721)

= 2 + 0.73926595

= 2.74 years

b-2. If you apply the discounted payback criterion, which investment will you choose?

Project B as it has lower  Lower Discounted Payback period.

c-1. What is the NPV for each project?

Project A

ear 0 1 2 3 4
Cashflow(in $)          (364,000)              46,000                     68,000                 68,000               458,000
PVF @11%                          1                 0.901                       0.812                   0.731                   0.659
Discounted Cashflow (Cash flow * PVF)    (364,000.00)        41,441.44               55,190.33           49,721.01         301,698.79

NPV = PV of Inflows - PV of Outflows

= (41441.44+55190.33+49721.01+301698.79) - 364000

= 448051.57-364000

= 84051.57

Project B

Year 0 1 2 3 4
Cashflow(in $)                (52,000)                 25,000                     22,000                  21,500                 17,500
PVF @11%                             1                   0.901                        0.812                    0.731                   0.659
Discounted Cashflow (Cash flow * PVF)                (52,000)                 22,522.52                     17,855.69                  15,720.61                 11,527.79

NPV = (22522.52+17855.69+15720.61+11527.79)-52000

= 67626.62-52000

= 15626.62

c-2. If you apply the NPV criterion, which investment will you choose?

Project A as it has Higher NPV

d-1. What is the IRR for each project?

Project A

IRR is the rate at which NPV=0. ie: PV of inflows = PV of outflows. It is calculated by trial and error method.

Lets find NPV at say 18%.

Year 0 1 2 3 4
Cashflow(in $)          (364,000)              46,000                     68,000                 68,000               458,000
PVF @18%                          1                 0.847                       0.718                   0.609                   0.516
Discounted Cashflow (Cash flow * PVF)    (364,000.00)        38,983.05               48,836.54           41,386.90         236,231.30

NPV1 = (38983.05+48836.54+41386.90+236231.30) - 364000

= 365437.80-364000

= 1437.80

Since NPV is positive, Take a higher rate say 19%

Year 0 1 2 3 4
Cashflow(in $)          (364,000)              46,000                     68,000                 68,000               458,000
PVF @19%                          1                 0.840                       0.706                   0.593                   0.499
Discounted Cashflow (Cash flow * PVF)    (364,000.00)        38,655.46               48,019.21           40,352.28         228,390.29

NPV2 = (38655.46+48019.21+40352.28+228390.29)-364000

= 355417.23-364000

= -8582.77

IRR = R1 + ((NPV1 * (R2 - R1)) / (NPV1 - NPV2))

= 18+((1437.80*(19-18)) / (1437.80+8582.77)

= 18+(1437.80/10020.57)

= 18+0.14348485166v

= 18.14%

Project B

Lets find NPV at say 25%

Year 0 1 2 3 4
Cashflow(in $)                (52,000)                 25,000                     22,000                  21,500                 17,500
PVF @25%                             1                   0.800                        0.640                    0.512                   0.410
Discounted Cashflow (Cash flow * PVF)          (52,000.00)           20,000.00               14,080.00            11,008.00             7,168.00

NPV = (20000+14080+11008+7168)-52000

= 52256-52000

= 256

Since NPV is positive, Take a higher rate say 26%

Year 0 1 2 3 4
Cashflow(in $)                (52,000)                 25,000                     22,000                  21,500                 17,500
PVF @26%                             1                   0.794                        0.630                    0.500                   0.397
Discounted Cashflow (Cash flow * PVF)          (52,000.00)           19,841.27               13,857.39            10,747.98             6,943.14

NPV = (19841.27+13857.39+10747.98+6943.14)-52000

= 51389.79-52000

= -610.22

IRR = R1 + ((NPV1 * (R2 - R1)) / (NPV1 - NPV2))

= 25 + ((256*(26-25)) / (256+610.22))

= 25 + (256/866.22)

= 25+0.29553693057

= 25.30%

d-2. If you apply the IRR criterion, which investment will you choose?

A project is acceptable when IRR is greater than required rate of return. Project B as this has higher IRR.

e-1. What is the profitability index for each project?

Project A

Profitability index = PV of future Cash Flows / Initial Investment

= 448051.57/364000

= 1.231

Project B

Profitability index = PV of future Cash Flows / Initial Investment

= 67626.62/52000

= 1.301

e-2. If you apply the profitability index criterion, which investment will you choose?

Project B as it has Higher Profitability index.

f. Based on your answers in (a) through (e), which project will you finally choose?

Project A as it has higher NPV as NPV supercedes all other evaluation technique.